INAF LBC+LUCIFER queue observing, 2010 January 14-15

Observer: V. Testa, F. Mannucci, D. Thompson
Telescope Operator: WWack
Instrument: LUCIFER
Data filename root: luci.20100115

Summary

  • 8 hours lost for snowing and high humidity
  • bad seeing during the remaining hours
  • 2 hours of scientific, on-source, possibly useful data with Lucifer
  • no problem with telescope and instrument, everything is working just fine

Data Summary

Files Project Object Description Notes
14-33 Turatto SN2009hv H, 30min seeing~1.7"
34-53 Turatto SN2009hv H, 30min seeing~1.4-1.9"
54-56 Turatto SN2009hv J, 5min seeing~2.5"
57-58 Turatto SN2009hv H, 3 min seeing~2.2"
60-79 Turatto SN2009hv H, 30 min seeing~2.3", thin clouds?
91-110 Turatto SN1009dc H, 30min seeing~1.5"

Details

01:00 - (6pm LT) Snowing
02:15 - Gettin better, high humidity
04:00 - Humidity 99%
06:00 - Humidity 100.7%
07:30 - Hymidity dropping
08:10 - Opening dome
08:13 - presetting, focusing (images 0001-0003)
08:34 - seeing: 3.5"in R; mirror is 4 deg. warmer (-6 to -2)
09:15 - seeing 3" in R
09:18 - SN2009hv - Turatto, H, 10sec x 9 x 20, 30min, turatto_2009hv_Hima.txt, (0014-0033), Alt~58
10:07 - SN2009hv again, 30min (0034-0053), seeing=1.9" in H
11:00 - SN2009hv, J, aborted due to bad seeing (2.5") (0054-0056)
11:14 - SN2009hv starting H again (0057-0058): aborted due to lost tracking (bad seeing, clouds?)
11:33 - SN2009hv starting again H (0060-0079) , seeing~2.3
12:23 - STD - BS9138 H band (0082-0086)
01:03 - SN2009dc - Turatto H band, 30 min (0091-0110) , possibly photometric
01:49 - STD - BS9160 H band (0111-0115)

-- FilippoMannucci - 15 Jan 2010
Topic revision: r9 - 16 Jan 2010, FilippoMannucci
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback